OpenStudy (daze96):

A man ran 20 miles in 2 hours. Calculate the average time that it took the man to run each mile in minutes

OpenStudy (amistre64):

how many minutes in 2 hours?

OpenStudy (daze96):

90

OpenStudy (amistre64):

60 + 60 = 90? try again

OpenStudy (anonymous):

120/20= 6 minutes per mile :)

OpenStudy (amistre64):

60+60 = how many minutes?

OpenStudy (daze96):

120

OpenStudy (amistre64):

good so the guy does: 20 miles in 120 minutes....lets divide these by 120 ------ ---------- 120 120 what do we get?

OpenStudy (amistre64):

120/120 = 1 so that 1 minute.... what is: 20 ---- ?? 120

OpenStudy (daze96):

ill be rite back

OpenStudy (daze96):

6

OpenStudy (gw2011):

2 hours = 120 minutes 120minutes/20 miles = 6 minutes/mile

OpenStudy (anonymous):

@amistre64: you are misleading them.. you are finding the amount of miles per minute, they want minutes per mile....

OpenStudy (amistre64):

20 2 ---- = --- = 1/6 120 12

OpenStudy (amistre64):

1/6 of a mile in 1 min

OpenStudy (amistre64):

misleading is a strong word lol...just read the problem wrong :)

OpenStudy (daze96):

wat im confused

OpenStudy (amistre64):

whats to be confused about

OpenStudy (daze96):

idk bt im just confused

OpenStudy (amistre64):

1/6 in 1 min; so he goes 1 mile in 6 mons

OpenStudy (amistre64):

6 mins

OpenStudy (daze96):

thats the answer

OpenStudy (amistre64):

aint misleading; just extra work to get to the right place lol

OpenStudy (daze96):

o becuz i have multi choice

OpenStudy (anonymous):

If you are still confused, this is what happened. I put the correct answer, amistre64 accidentally misread the question and helped you solve the inverse of the correct answer.

OpenStudy (amistre64):

:)

OpenStudy (daze96):

o

OpenStudy (daze96):

thnx